# R1-Code-Interpreter: LLMs Reason with Code via Supervised and Multi-stage Reinforcement Learning

### v2

@@ 1 Introduction
While reinforcement learning (RL)-based fine-tuning has significantly improved LLMs’ reasoning and planning Wang et al. (2024); Guo et al. (2025); Jaech et al. (2024), models still struggle with seemingly simple tasks (Chen et al., 2025) and incur high token costs during inference-time search (Chen et al., 2024a). Textual reasoning excels at semantics and commonsense, but falls short in precise computation, symbolic manipulation, and algorithmic processing (Valmeekam et al., 2022). In contrast, symbolic code generation handles these rigorously and benefits from external tools (e.g., equation solvers). Prompting LLMs to generate and execute code often outperforms pure textual reasoning (Madaan et al., 2022; Liang et al., 2022; Chen et al., 2022).
@@ 2 Challenges in 144 Reasoning and Planning Tasks
We compile 144 tasks from three major reasoning and planning benchmarks: 33 from SymBench (Chen et al., 2025), 27 from Big-Bench-Hard (Suzgun et al., 2022), and 84 from Reasoning-Gym [^1] https://github.com/open-thought/reasoning-gym. After removing near-duplicates, each task retains over 200 diverse samples. All tasks are standardized into a unified format and evaluated using rule-based criteria (e.g., exact match or constraint checks) for efficient rollout and testing. The tasks cover diverse reasoning and planning challenges for LLM evaluation. Detailed task descriptions are in Appendix Sec. E, and their associated capability categories, logic, spatial, order, optimization, search, and math, are summarized in Appendix Table LABEL:tab:SymBench_class. Recent work Chen et al. (2024c) highlights key challenges in steering LLMs to choose effectively between code and text: (1) The better mode (code vs. text) depends on task type, complexity, and model capabilities, which is often hard to predict; (2) LLM-generated code frequently degenerates into hard-coded text-like scripts, limiting its symbolic utility (Yang et al., 2024); (3) Many challenging tasks require models to iteratively explore and optimize the solutions in a difficult multi-turn behavior. Appendix Fig. 7 shows a representative example of GPT-5 failing to use the Code Interpreter.
@@ 4.2 Bottlenecks in GRPO training of Code Interpreter
Hence the policy signal is maximized at $p=\tfrac{1}{2}$ and vanishes as $p\to 0$ or $1$. In batches dominated by too-easy or too-hard items, the update is therefore governed by the KL term, contracting $\pi_{\theta}$ toward $\pi_{\text{ref}}$. Hence, optimization cannot make headway. See Appendix Sec. C for detailed proving and articulation.
@@ 4.3 Multi-stage curriculum learning with potential measurement
Multi-stage curriculum: To implement curriculum learning, we sort all training samples by $\Pi_{i}$ and partition them into four equally sized groups, from highest to lowest potential. The partition is sample-wise rather than task-wise, as samples within the same task can differ significantly in potential due to varying difficulty. GRPO training begins with the highest-potential group, then progressively incorporates lower-potential groups in subsequent stages. Each stage runs for 150 steps, gradually expanding the training distribution until the full dataset is included by stage 4. Because the usable gradient scales with p(1p)p(1-p), ranking by $\Pi_{i}$ prioritizes samples with maximal expected signal. Fig. 5 shows training rewards and test scores over steps. Contrary to training without curriculum learning (Fig. 3), both metrics rise apparently, particularly in the first two stages. We also observe sharp drops in reward when new samples are merged at each new stage. However, little improvement is gained in the final stage, suggesting that adding low-potential samples offers limited benefit. These observations align with our theoretical analysis.
@@
Full Name Description (15 agents)
@@ 5.3 Ablation study
Why multi-turn Code Interpreter? An important question is how much multi-turn code execution contributes to improving LLM reasoning, i.e., whether the observed gains primarily stem from in-domain training that other frameworks could also achieve. To investigate this, we train the LLM with All Text and All Code using the same amount of GPT-4o sampled training data, as shown in Appendix Fig. 9. Both All Text and All Code yield smaller improvements than CI, particularly on testing tasks, indicating that the multi-turn CI framework is more generalizable and effective, and that the gains are not primarily due to in-domain training.
@@ 5.4 R1-CI training and inference cost
6 Related Work
@@
7 Conclusion
Appendix A Example of GPT-5 failure in code/text decision
@@
Appendix B Sparse GPU utilization due to time-consuming code execution
@@
Appendix C Why vanilla GRPO fails in mixed difficulty, and why potential-guided curriculum works
C.1 Preliminaries and notation
@@
$\nabla_{\theta}\mathcal{J}_{\mathrm{GRPO}}(\theta)=\underbrace{\frac{1}{G}\sum_{i=1}^{G}(r_{i}-\bar{r})\,v_{i}}_{\text{policy term}}\;-\;\underbrace{\beta\,\nabla_{\theta}D_{\mathrm{KL}}\!\bigl(\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\bigr)}_{\text{regularizer}},\qquad v_{i}\;:=\;\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\nabla_{\theta}\log\pi_{\theta}(y_{i,t}\mid x,y_{i,<t};\mathcal{C}).$ (C.1)
@@
C.2 A variance identity for group-relative Bernoulli rewards
Lemma C.1 (Within-group Bernoulli variance).
@@ C.2 A variance identity for group-relative Bernoulli rewards
$\mathbb{E}\big[(r_{i}-\bar{r})^{2}\big]=\mathrm{Var}(r_{i}-\bar{r})=p(1-p)\!\left(1-\frac{1}{G}\right).$ (C.2)
@@
C.3 A bound on the policy-gradient magnitude
@@
Let $g_{\mathrm{pol}}:=\tfrac{1}{G}\sum_{i=1}^{G}(r_{i}-\bar{r})\,v_{i}$ be the policy term in equation C.1. Then
$\mathbb{E}\big[\|g_{\mathrm{pol}}\|^{2}\big]\;\leq\;p(1-p)\!\left(1-\frac{1}{G}\right)\;\mathbb{E}\!\left[\frac{1}{G}\sum_{i=1}^{G}\|v_{i}\|^{2}\right].$ (C.3)
@@
With $a_{i}=r_{i}-\bar{r}$ and the prefactor $1/G$, we get $\|g_{\mathrm{pol}}\|^{2}\leq\frac{1}{G^{2}}\Big(\sum_{i}(r_{i}-\bar{r})^{2}\Big)\Big(\sum_{i}\|v_{i}\|^{2}\Big)$. Taking expectation and using Lemma C.1, $\mathbb{E}\big[\sum_{i}(r_{i}-\bar{r})^{2}\big]=G\,p(1-p)\!\left(1-\tfrac{1}{G}\right)$, yielding equation C.3. ∎
Corollary C.1 (Vanishing signal at the extremes).
The upper bound in equation C.3 vanishes as $p\to 0$ or $p\to 1$, and is maximized at $p=\tfrac{1}{2}$.
C.4 Rebuttal of “given enough steps, GRPO will succeed anyway”
In a heterogeneous batch dominated by too-easy or too-hard items, $p(1-p)\approx 0$ for most samples, so equation C.3 implies
$\mathbb{E}\big[\|g_{\mathrm{pol}}\|^{2}\big]\;\approx\;0.$ (C.4)
Consequently, the update in equation C.1 is governed by the regularizer $\beta\,\nabla_{\theta}D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}})$, which contracts $\pi_{\theta}$ back toward $\pi_{\mathrm{ref}}$. No amount of additional steps recovers signal from items with identically zero Bernoulli variance.
C.5 Descent lemma view: expected improvement per step
@@
$\mathbb{E}\big[\mathcal{J}_{\mathrm{GRPO}}(\theta^{+})-\mathcal{J}_{\mathrm{GRPO}}(\theta)\big]\;\geq\;\eta\,\big\|\mathbb{E}[\widehat{g}]\big\|^{2}\;-\;\frac{L\eta^{2}}{2}\,\mathbb{E}\big[\|\widehat{g}\|^{2}\big].$ (C.5)
Potential-guided sampling (defined below) increases $\big\|\mathbb{E}[\widehat{g}_{\mathrm{pol}}]\big\|$ and decreases the fraction of near-zero-variance items, thereby increasing the first term in equation C.5 at fixed batch size. In contrast, vanilla GRPO on mixed difficulty often drives $\big\|\mathbb{E}[\widehat{g}_{\mathrm{pol}}]\big\|\!\downarrow$ toward zero by equation C.3, making the KL term dominate the dynamics.
C.6 Potential as a proxy for learning signal
@@
$\hat{p}_{i}\;:=\;\frac{1}{N}\sum_{j=1}^{N}\mathbf{1}\{\text{answer }j\text{ is correct}\},\qquad\Pi_{i}\;:=\;4\,\hat{p}_{i}\,(1-\hat{p}_{i})\in[0,1].$ (C.6)
Lemma C.2 (Concentration of the potential estimator).
@@
Let $v_{i}$ be defined as in equation C.1. Up to the slowly-varying factor $\mathbb{E}[\|v_{i}\|^{2}]$, the bound in equation C.3 shows that $\mathbb{E}\big[\|g_{\mathrm{pol}}\|^{2}\big]$ is maximized when $p_{i}(1-p_{i})$ is maximized, i.e. near $p_{i}=\tfrac{1}{2}$. Therefore $\Pi_{i}$ serves as a proxy for per-item learning signal.
C.7 Why multi-stage curriculum helps
Let $\mathcal{B}$ denote the distribution over batch items. Define $\psi(\mathcal{B}):=\mathbb{E}_{x\sim\mathcal{B}}[\,p_{x}(1-p_{x})\,].$ From equation C.3, larger $\psi(\mathcal{B})$ increases the policy term’s expected squared norm. A potential-ranked curriculum selects items with large $\Pi_{i}$ in early stages and gradually lowers the threshold, so that previously too-hard items (with small $p_{i}$) enter the curriculum when training has moved them into the high-variance region. Thus the curriculum maintains batches with large $\psi(\mathcal{B})$ throughout training, improving the expected gain equation C.5.
Remark C.1 (Clipping and KL).
Our analysis ignores clipping in equation C.1; including it only reduces the policy term’s magnitude, strengthening the conclusion that mixed-difficulty batches yield weak signal. The KL regularizer is essential for stability but dominates updates exactly when the policy signal is small.
C.8 Conclusions

### v3

@@ 1 Introduction
While reinforcement learning (RL)-based fine-tuning has significantly improved LLMs’ reasoning and planning Wang et al. (2024); Guo et al. (2025); Jaech et al. (2024), models still struggle with seemingly simple tasks (Chen et al., 2025) and incur high token costs during inference-time search (Chen et al., 2024a). Textual reasoning excels at semantics and commonsense, but falls short in precise computation, symbolic manipulation, optimization, and algorithmic processing (Valmeekam et al., 2022). In contrast, symbolic code generation handles these rigorously and benefits from external tools (e.g., equation solvers). Prompting LLMs to generate and execute code often outperforms pure textual reasoning (Madaan et al., 2022; Liang et al., 2022; Chen et al., 2022).
@@ 2 Challenges in 144 Reasoning and Planning Tasks
We compile 144 tasks from three major reasoning and planning benchmarks: 33 from SymBench (Chen et al., 2025), 27 from Big-Bench-Hard (Suzgun et al., 2022), and 84 from Reasoning-Gym [^1] https://github.com/open-thought/reasoning-gym. After removing near-duplicates, each task retains over 200 diverse samples. All tasks are standardized into a unified format and evaluated using rule-based criteria (e.g., exact match or constraint checks) for efficient rollout and testing. The tasks cover diverse reasoning and planning challenges for LLM evaluation. Detailed task descriptions are in Appendix Sec. H, and their associated capability categories, logic, spatial, order, optimization, search, and math, are summarized in Appendix Table LABEL:tab:SymBench_class. Recent work Chen et al. (2024c) highlights key challenges in steering LLMs to choose effectively between code and text: (1) The better mode (code vs. text) depends on task type, complexity, and model capabilities, which is often hard to predict; (2) LLM-generated code frequently degenerates into hard-coded text-like scripts, limiting its symbolic utility (Yang et al., 2024); (3) Many challenging tasks require models to iteratively explore and optimize the solutions in a difficult multi-turn behavior. Appendix Fig. 11 shows a representative example of GPT-5 failing to use the Code Interpreter.
@@ 4.2 Bottlenecks in GRPO training of Code Interpreter
Hence the policy signal is maximized at $p=\tfrac{1}{2}$ and vanishes as $p\to 0$ or $1$. In batches dominated by too-easy or too-hard items, the update is therefore governed by the KL term, contracting $\pi_{\theta}$ toward $\pi_{\text{ref}}$. Hence, optimization cannot make headway. See Appendix Sec. D for detailed proving and articulation.
@@ 4.3 Multi-stage curriculum learning with potential measurement
Multi-stage curriculum: To implement curriculum learning, we sort all training samples by $\Pi_{i}$ and partition them into four equally sized groups (four group IP ranges are: [0.0, 0.32], [0.32, 0.48], [0.48, 0.64], and [0.64, 1.00]), from highest to lowest potential. The partition is sample-wise rather than task-wise, as samples within the same task can differ significantly in potential due to varying difficulty. GRPO training begins with the highest-potential group, then progressively incorporates lower-potential groups in subsequent stages. Each stage runs for 150 steps, gradually expanding the training distribution until the full dataset is included by stage 4. Because the usable gradient scales with p(1p)p(1-p), ranking by $\Pi_{i}$ prioritizes samples with maximal expected signal. Fig. 5 shows training rewards and test scores over steps. Contrary to training without curriculum learning (Fig. 3), both metrics rise apparently, particularly in the first two stages. We also observe sharp drops in reward when new samples are merged at each new stage. However, little improvement is gained in the final stage, suggesting that adding low-potential samples offers limited benefit. These observations align with our theoretical analysis.
@@
Full Name Description (4 agents)
@@ 5.1 Overall Better Performance
Evaluation on out-of-distribution (OOD) tasks. In Appendix Table 4, we evaluate the performance of R1-CI-7B and R1-CI-14B on unseen OOD tasks, including Graduate-Level Google-Proof Q&A (GPQA, Diamond) (Rein et al., 2024) and the American Invitational Mathematics Examination (AIME 24&25). Both models significantly outperform their untrained counterparts. In Appendix Fig. 13, we further assess the generalizability of our training framework by training the 14B model with SymBench and Reasoning-Gym as training data, and evaluating it on BBH as an OOD benchmark. The number of training tasks remained 107 as before. The green curve in the figure shows the new model’s performance on BBH, which is comparable to the purple curve representing the original R1-CI-14B model. These results demonstrate the satisfying generalizability of R1-Code-Interpreter to unseen tasks from diverse sources.
@@ 5.2 Response characteristics
Figure 7: Comparison of GRPO, PPO, and Reinforce++ as RL algorithms.
Figure 8: Offline vs. online RL for multi-stage curriculum learning.
Figure 9: Ablation studies on SFT dataset design and training frameworks. All experiments are trained from 7B models. (a) We compare SFT performance across three frameworks: multi-turn Code Interpreter (CI), single-turn All Text, and single-turn All Code. (b) We evaluate alternative SFT dataset designs: (1) including incorrect answers, (2) omitting prompt variation during synthesis, and (3) removing the multi-turn interaction emphasis.
Figure 10: Continued multi-stage curriculum learning. We conduct a second stage GRPO training, incorporating updated improvement potentials and a newly partitioned dataset.
@@ 5.3 Ablation study
Multi-turn Code Interpreter: To investigate the gains from the multi-turn framework, we train the LLM with All Text and All Code using the same amount (6.5k) of GPT-4o sampled training data, as shown in Fig. 10a. Both All Text and All Code yield smaller improvements than CI, particularly on testing tasks, indicating that the multi-turn CI framework is more generalizable and effective, and that the gains are not primarily due to in-domain training.
SFT dataset design In Fig. 10b, we further evaluate alternative SFT dataset designs: (1) w/ Wrong Data: In addition to the existing 6.5k correct answer trajectories, we also keep the same-sized incorrect answers synthesized by GPT-4o. However, this setting degrades model performance and increases variance, indicating instability. (2) w/o Varied Prompts: We synthesize the SFT dataset using a single generic prompt without varying code-use strategies. Results show that diverse prompts and strategies are crucial for enhancing dataset diversity and overall performance. (3) w/o Multi-Turn Emphasis: We remove the emphasis on multi-turn answer trajectories and adaptive solving strategies. This ablation leads to noticeable performance drops, demonstrating that diverse and high-quality multi-turn data play a key role in improving model capability.
@@
GRPO vs. PPO vs. Reinforce++:  In Fig. 8, we evaluate the effectiveness of our framework with alternative RL algorithms. We compare GRPO (Shao et al., 2024), PPO (Schulman et al., 2017), and Reinforce++ (Hu, 2025), and observe that their performances on both training rewards and test scores are comparable. The overall upward performance trend indicates that our framework is adaptable across different RL algorithms.
Offline vs. online RL training:  In our multi-stage RL training setup, the dataset is partitioned into four groups before the first stage and remains fixed throughout, corresponding to an offline training setting. In Fig. 8, we compare this with two online variants: (1) Online-Merge: Before each stage, the dataset is repartitioned using the current model, and training proceeds on these updated groups. As in the offline case, later stages merge data from higher to lower potential groups, so the final stage accesses all data. (2) Online-Unmerge: Similar to (1), but each stage trains only on a single unmerged group, proceeding from high to low potential groups across stages. We find that offline and online training achieve comparable performance in later stages. Although the online variant converges faster, it incurs higher inference costs due to repeated repartitioning. The unmerged online setting performs worse overall. These results suggest that offline training is comparable to the relatively costly online training in our setting.
RL training round number:  In Fig. 10, we extend the RL training process with a newly partitioned dataset, derived from the first-round model shown in Fig. 5. However, the testing score does not improve with continued training; in fact, during the first stage, the testing score decreases even as training rewards increase. We attribute this to overfitting on the limited data group used in stage 1, which is partially alleviated in later stages as more data are introduced. These results indicate that a single round of RL training is sufficient, as the model effectively adapts across the entire dataset.
Impact of sample improvement potential on GRPO training:  In Fig. 14, we evaluate the effect of training GRPO with datasets of equal size but different ranges of improvement potential. Training samples are drawn from Reasoning Gym and SymBench, while evaluation is conducted on BBH for fair comparison. Models trained on higher-potential samples $[0.64,1.00]$ exhibit a more pronounced increase in training rewards and achieve higher BBH test scores. These results support our theoretical analysis in Section 4.3 and validate the correctness of multi-stage curriculum learning.
@@ 5.4 R1-CI training and inference cost
6 Conclusion
7 Ethics Statement
This paper contributes to advancing Foundation Models by augmenting language models with a Code Interpreter, which has strong potential to improve safety, performance, and alignment with human preferences. However, such capabilities are inherently dual-use, the same techniques that augment models toward harmless outputs can, with minor changes, be misused to generate harmful content. While misuse is a concern, we believe the broader societal benefits outweigh the risks.
8 Reproducibility Statement
For better reproducibility, we include detailed descriptions of our curated 144 tasks in Appendix Sec. H and the synthesis of SFT and GRPO training datasets in Sec. 4. The theoretical analysis of multi-stage GRPO training and the measurement of improvement potential are included in Appendix Sec. D, Sec. 4.2, and Sec. 4.3. The full code of training and dataset synthesis are attached in Supplementary Material of the submission. Our code, model, and dataset will be made publicly available under an open-source license following the acceptance of the paper.
9 Large Language Model Usage for Writing
In this paper, we use LLMs—specifically ChatGPT—as general-purpose writing aids. Draft text was provided to these models for grammatical correction and structural refinement, after which the output was verified and further edited when necessary. Their use was strictly limited to text refinement; they were not employed to generate new content or references.
10 Acknowledgments
This work was supported by ONR under Award N00014-22-1-2478 and MIT-IBM Watson AI Lab. However, this article solely reflects the opinions and conclusions of its authors.
Appendix A Related Work
@@
Appendix B Example of GPT-5 failure in code/text decision
@@
Appendix C Sparse GPU utilization due to time-consuming code execution
@@
Appendix D Why vanilla GRPO fails in mixed difficulty, and why potential-guided curriculum works
D.1 Preliminaries and notation
@@
$\nabla_{\theta}\mathcal{J}_{\mathrm{GRPO}}(\theta)=\underbrace{\frac{1}{G}\sum_{i=1}^{G}(r_{i}-\bar{r})\,v_{i}}_{\text{policy term}}\;-\;\underbrace{\beta\,\nabla_{\theta}D_{\mathrm{KL}}\!\bigl(\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\bigr)}_{\text{regularizer}},\qquad v_{i}\;:=\;\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\nabla_{\theta}\log\pi_{\theta}(y_{i,t}\mid x,y_{i,<t};\mathcal{C}).$ (D.1)
@@
D.2 A variance identity for group-relative Bernoulli rewards
Lemma D.1 (Within-group Bernoulli variance).
@@ D.2 A variance identity for group-relative Bernoulli rewards
$\mathbb{E}\big[(r_{i}-\bar{r})^{2}\big]=\mathrm{Var}(r_{i}-\bar{r})=p(1-p)\!\left(1-\frac{1}{G}\right).$ (D.2)
@@
D.3 A bound on the policy-gradient magnitude
@@
Let $g_{\mathrm{pol}}:=\tfrac{1}{G}\sum_{i=1}^{G}(r_{i}-\bar{r})\,v_{i}$ be the policy term in equation D.1. Then
$\mathbb{E}\big[\|g_{\mathrm{pol}}\|^{2}\big]\;\leq\;p(1-p)\!\left(1-\frac{1}{G}\right)\;\mathbb{E}\!\left[\frac{1}{G}\sum_{i=1}^{G}\|v_{i}\|^{2}\right].$ (D.3)
@@
With $a_{i}=r_{i}-\bar{r}$ and the prefactor $1/G$, we get $\|g_{\mathrm{pol}}\|^{2}\leq\frac{1}{G^{2}}\Big(\sum_{i}(r_{i}-\bar{r})^{2}\Big)\Big(\sum_{i}\|v_{i}\|^{2}\Big)$. Taking expectation and using Lemma D.1, $\mathbb{E}\big[\sum_{i}(r_{i}-\bar{r})^{2}\big]=G\,p(1-p)\!\left(1-\tfrac{1}{G}\right)$, yielding equation D.3. ∎
Corollary D.1 (Vanishing signal at the extremes).
The upper bound in equation D.3 vanishes as $p\to 0$ or $p\to 1$, and is maximized at $p=\tfrac{1}{2}$.
D.4 Rebuttal of “given enough steps, GRPO will succeed anyway”
In a heterogeneous batch dominated by too-easy or too-hard items, $p(1-p)\approx 0$ for most samples, so equation D.3 implies
$\mathbb{E}\big[\|g_{\mathrm{pol}}\|^{2}\big]\;\approx\;0.$ (D.4)
Consequently, the update in equation D.1 is governed by the regularizer $\beta\,\nabla_{\theta}D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}})$, which contracts $\pi_{\theta}$ back toward $\pi_{\mathrm{ref}}$. No amount of additional steps recovers signal from items with identically zero Bernoulli variance.
D.5 Descent lemma view: expected improvement per step
@@
$\mathbb{E}\big[\mathcal{J}_{\mathrm{GRPO}}(\theta^{+})-\mathcal{J}_{\mathrm{GRPO}}(\theta)\big]\;\geq\;\eta\,\big\|\mathbb{E}[\widehat{g}]\big\|^{2}\;-\;\frac{L\eta^{2}}{2}\,\mathbb{E}\big[\|\widehat{g}\|^{2}\big].$ (D.5)
Potential-guided sampling (defined below) increases $\big\|\mathbb{E}[\widehat{g}_{\mathrm{pol}}]\big\|$ and decreases the fraction of near-zero-variance items, thereby increasing the first term in equation D.5 at fixed batch size. In contrast, vanilla GRPO on mixed difficulty often drives $\big\|\mathbb{E}[\widehat{g}_{\mathrm{pol}}]\big\|\!\downarrow$ toward zero by equation D.3, making the KL term dominate the dynamics.
D.6 Potential as a proxy for learning signal
@@
$\hat{p}_{i}\;:=\;\frac{1}{N}\sum_{j=1}^{N}\mathbf{1}\{\text{answer }j\text{ is correct}\},\qquad\Pi_{i}\;:=\;4\,\hat{p}_{i}\,(1-\hat{p}_{i})\in[0,1].$ (D.6)
Lemma D.2 (Concentration of the potential estimator).
@@
Let $v_{i}$ be defined as in equation D.1. Up to the slowly-varying factor $\mathbb{E}[\|v_{i}\|^{2}]$, the bound in equation D.3 shows that $\mathbb{E}\big[\|g_{\mathrm{pol}}\|^{2}\big]$ is maximized when $p_{i}(1-p_{i})$ is maximized, i.e. near $p_{i}=\tfrac{1}{2}$. Therefore $\Pi_{i}$ serves as a proxy for per-item learning signal.
D.7 Why multi-stage curriculum helps
Let $\mathcal{B}$ denote the distribution over batch items. Define $\psi(\mathcal{B}):=\mathbb{E}_{x\sim\mathcal{B}}[\,p_{x}(1-p_{x})\,].$ From equation D.3, larger $\psi(\mathcal{B})$ increases the policy term’s expected squared norm. A potential-ranked curriculum selects items with large $\Pi_{i}$ in early stages and gradually lowers the threshold, so that previously too-hard items (with small $p_{i}$) enter the curriculum when training has moved them into the high-variance region. Thus the curriculum maintains batches with large $\psi(\mathcal{B})$ throughout training, improving the expected gain equation D.5.
Remark D.1 (Clipping and KL).
Our analysis ignores clipping in equation D.1; including it only reduces the policy term’s magnitude, strengthening the conclusion that mixed-difficulty batches yield weak signal. The KL regularizer is essential for stability but dominates updates exactly when the policy signal is small.
D.8 Conclusions